OpenStudy (anonymous):

Solve x2 + 5x = -6. x = 5 and x = 1 x = -5 and x = -1 x = 3 and x = 2 x = -3 and x = -2

OpenStudy (nurali):

x2+5x=-6 x2+5x+6=0 x2+3x+2x+6=0 x(x+3)+2(x+3)=0 (x+3)(x+2)=0 x=-3 and x=-2

OpenStudy (anonymous):

thanks again! love you(: